Consider two stocks. Stock A has a standard deviation of 56% and stock B has a standard deviation of 37%. The stocks have a correlation of -0.05. You plan to invest $6,113 into stock A and $6,720 into stock B. What is the standard deviation of your two stock portfolio? (round weights to 3 decimal places and final answer to 2 decimal places).
Answer #1

Standard deviation of portfolio:

Oportfolio = wig+wo3 + 2wW2P1,20102 Where: W W 04 02 P12 = = = = = Proportion of the portfolio invested in Asset 1 Proportion

w1 = $6,113/($6,113 + $6,720) = 0.476

w2 = $6,720/($6,113 + $6,720) = 0.524

o= V(0.476 * 0.56)2 + (0.524 * 0.37)2 + (2*0.476 * 0.524 * 0.56*0.37 * -0,05)

0 = V0.071159 +0.037539 -0.00517

0 = V0.10353

0 =0.3218

Standard deviation = 32.18%